OpenStudy (anonymous):

Am I right? Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (0, 4) and (-2, 8) (3, 7) and (-1, -5) The slope of Set #1 is smaller than the slope of Set #2.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Slope of Set #1: \[\Large \frac{{8 - 4}}{{ - 2 - 0}} = \frac{4}{{ - 2}} = - 2\] Slope of Set #2: \[\Large \frac{{ - 5 - 7}}{{ - 1 - 3}} = \frac{{ - 12}}{{ - 4}} = 3\]

OpenStudy (anonymous):

Yay!! :DD I got it right, then! :DDD

hero (hero):

The Hero, if you're going to offer an answer for us to check, you should include all of the work you did as well.

OpenStudy (anonymous):

Umm, Alright then...